# TREACLE: Thrifty Reasoning via Context-Aware LLM and Prompt Selection

### v1

@@ 1 Introduction
The success of large language models (LLMs) in recent years has led to a explosion of heterogeneous models and providers, including as Meta’s Llama, OpenAI’s ChatGPT, and Google’s Gemini. As LLMs continue to proliferate in the near future, we envisage a generative AI marketplace with a large variety of providers, LLMs, and deployments. Notably, LLMs have widely varying capabilities and costs: capabilities in terms of accuracy in responding to different types of queries, and cost in terms of monetary price and query latency. As an illustration, the accuracy versus cost tradeoffs of various Llama and GPT LLMs are shown in Fig. 1 on grade school math word problems (Cobbe et al., 2021). As can be seen, GPT-3.5 tends to have lower accuracy than GPT-4 (79% vs 92% respectively), but costs about 20 times less. This heterogeneous array of LLMs can bewilder users who must choose between them.
Figure 1: ${\sf\small TREACLE}$ chooses between LLMs to achieve high accuracy and $\sim$85cost reduction compared to individual LLMs.
Another challenge is that the specific prompt included in the question plays a critical role in eliciting accurate responses. This is especially true for reasoning problems where prompting a model to explain its reasoning can produce more accurate, but often more costly, answers. Chain-of-thought (CoT) (Wei et al., 2022) is an example of such a prompting scheme, in which the question includes a few examples of worked out problems, which cost more (due to the additional words included in the question) but also produce more accurate responses. For example, in Figure 1, GPT-4 with CoT (pink triangle) achieves a 92% accuracy, compared to GPT-4 with a domain expert prompt (brown dot, reminding the LLM that it is a “math solver”) that achieves 83%. However, using the CoT prompt costs 3.9$\times$ more due to the extra words included in the query. A final challenge is that the optimal choice of LLM and prompt depends on the specific question being asked; the accuracy of a particular LLM and prompt combination for a particular question is unknown in advance, requiring learning or prediction.
@@
Several recent works utilize multiple LLMs during inference with a focus on cascade design and accuracy-cost tradeoffs. Most aim to maximize accuracy and lack an explicit way to control long-term costs, as TREACLE has. More broadly, by posing the problem of LLM and prompt selection as a budget-constrained policy optimization, TREACLE provides a unified approach to efficient LLM cascades (see Table 1). TREACLE’s policy can make informed decisions based on the full context of the LLM cascade, including the query embedding, answer statistics, and remaining budget.
Table 1: Comparison to related works. TREACLE poses the problem of LLM and prompt selection as a policy optimization. This policy seamlessly admits query embedding, long term budget, output statistics/consistency, and more as its input.
Query embedding Consistency Prompt and LLM selection Long term budget Robust to new models
FrugalGPT (Chen et al., 2023)
AutoMix (Madaan et al., 2023)
Mixture of Thought (Yue et al., 2023)
TREACLE (ours)
Overall, this paper makes the following contributions:
@@
  • An adaptive LLM and prompt selection policy based on reinforcement learning. ${\sf\small TREACLE}$ dynamically chooses the right LLM and prompt for each question. It does this by leveraging context about the current question, re-querying the models if needed to verify the consistency of the responses, and thinking ahead about the remaining budget.
  • Extensive evaluations. We show that TREACLE substantially saves on cost while maintaining high accuracy on challenging mathematical and commonsense reasoning tasks. We demonstrate its robustness to different budgets, time-varying costs, question difficulty, price changes, new LLMs, and new unseen task types.
@@ 2 Related Work
FrugalGPT (Chen et al., 2023) is perhaps the closest to this work, as they considered a similar cost-constrained LLM selection problem with a threshold-based policy to select from a sorted list of LLMs. Our approach differs in several key aspects: we utilize a reinforcement learning policy that chooses both LLMs and prompts, rather than a threshold-based scheme; we utilize the full context of the current question to make decisions, including the text embedding of the current question and the history of past responses; and our method can re-query the same LLM and aggregate the responses of previously queried LLMs to estimate the correctness of the current response.
Mixture of Thought (Yue et al., 2023) explored the idea of response consistency in order to choose the right LLMs. The intuition is that higher consistency in the re-queries implies higher confidence in the correctness of the response. TREACLE employs response consistency as an input feature, along with other features, for LLM selection. AutoMix (Madaan et al., 2023) introduces a “meta-verifier” to estimate whether a response is correct or a more powerful LLM is needed. Both works measure cost as a by-product of combining multiple LLMs rather than as a core long-term constraint across an entire set of questions, as we do.
Other lines of work include uncertainty estimation or prompt engineering to improve accuracy (Lin et al., 2022; Xiong et al., 2023; Yue et al., 2023; Si et al., 2023; Cai et al., 2023; Naik et al., 2023), which is complementary to our work. The related work is summarized in Table 1.
@@ 3 Problem Statement
We study the standard natural language query problem of providing correct responses to a series of questions. We focus on reasoning problems (e.g., grade school math problems) because they are challenging with multiple logical steps required to reach a final correct response. The problem involves answering a sequence of $n$ questions $\mathcal{Q}$ with correct responses $\mathcal{Y}$; in other words, we have a set of questions and responses $\{(Q_{1},Y_{1}),(Q_{2},Y_{2}),\ldots,(Q_{n},Y_{n})\}$. We have a set $\mathcal{M}$ of $m$ large language models (LLMs) at our disposal, which can be accessed either locally or remotely through APIs: =$\mathcal{M}=\{M_{1},M_{2},\ldots,M_{m}\}$. Also, we have a choice between $p$ prompt types, $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{p}\}$. These models and prompts have different costs (in terms of latency and monetary price) and accuracy.
The primary goal is to ensure that as many responses as possible are correct, while simultaneously minimizing the associated costs. Thus, we seek to maximize the overall performance $\mathbb{E}_{(Q,Y)}\texttt{reward}(Y,\hat{Y})$ while ensuring the total cost is bounded by a user-defined budget $B$, i.e., $\sum_{i=1}^{n}\texttt{cost}(Q_{i})\leq B$. Here, $\texttt{cost}(\cdot)$ is the cost of processing a question, in terms of latency and monetary price (details below). The $\texttt{reward}(\cdot)$ function measures the correctness of the generated response (i.e., the accuracy). $\hat{Y}$ is the response returned by TREACLE.
@@ 4 Proposed Framework: TREACLE
To solve this problem, we propose the TREACLE framework, depicted in Figure 2. Let the possible combinations of language models and prompts be denoted by $\{(M+P)_{1},(M+P)_{2},\ldots,(M+P)_{K}\}$, where $K\leq mp$. When a new question $Q_{i}$ arrives, TREACLE starts by selecting a model $M\in\mathcal{M}$ and choosing an associated prompt $P\in\mathcal{P}$ to generate a response, denoted as $\hat{O}_{i}=M(Q_{i},P)$. TREACLE returns this as the final response $\hat{Y}_{i}$ for this question if it has a high degree of confidence in its correctness, and deducts the cost of the question and its response from the total budget $B$. Otherwise, TREACLE can select another LLM $M$ and prompt $P$ (whose choice may be informed by the result of all previously chosen models, prompts, and their responses) and re-query. This iterative process continues until TREACLE returns a final response (based on its learned policy). TREACLE then proceeds to the next question with the remaining budget and repeats the process, until all questions have been answered or there is no remaining budget. We use the terminology that one “question” may have multiple “re-queries” to different LLMs.
Figure 2: Overview of TREACLE framework. TREACLE decides on the next (LLM, prompt) pair to query in a context-aware fashion which is summarized in the State variable.
We model the problem as a Markov decision process, consisting of a set of states $\mathcal{S}$ and a set of actions $\mathcal{A}$. When a question arrives, the process is in a certain state $s$, and the decision maker may choose any action $a$ that is available in that state $s$. After an action is taken, the process responds by moving to a new state $s^{\prime}$, giving the decision maker a corresponding reward $R_{a}\left(s,s^{\prime}\right)$.
@@
  • Response consistency: Records all previous responses and the normalized frequency of their occurrences. The intuition is that the consistency of the previous responses can be used as a measure of confidence in the response correctness (Wang et al., 2022; Madaan et al., 2023).
@@
  • Current question’s text embedding: Intuitively, we want to capture the question type or difficulty, which can impact the model and prompt selection decision. TREACLE does this using a text embedding of the query (Greene et al., 2022).
@@
  • Normalized remaining budget: Based on the remaining budget, we compute the estimated number of queries for each model prompt pair as follows: $\mathcal{B}_{k}=\frac{\text{total remaining budget}}{(\text{\# questions remaining})(\text{avg cost per query of $(M+P)_{k}$})}$. The average cost per query is estimated based on the questions seen so far. If there is a large remaining budget, TREACLE may consider re-querying with large models, but when there is less remaining budget, the policy may wish to use a cheaper model or a terse prompt.
@@
By allowing re-querying (action $a_{2}$), the current action influences the next state, by impacting the question under consideration and thus the relevant state features, making this a non-trivial MDP. For $a_{3}$, we constrained the set of possible model-prompt pairs to a sorted list, which we experimentally found worked better. In other words, instead of allowing TREACLE to select any possible model and prompting scheme, we sort the $(M+P)_{k}$ in ascending order of model accuracy and cost and only allow TREACLE to select the next element in this list $(M+P)_{k+1}$. The ordering is determined by our experimental characterization of different $(M+P)$ combinations (§5.1.2) with theoretical justification in Appendix B.
@@
Rewards. The reward function assigns a positive reward to correct responses. Specifically, $R_{a}\left(s,s^{\prime}\right)=\mathbb{P}\left[\hat{Y}=Y|a=a_{1}\right]+\lambda\mathbb{P}\left[\hat{O}=Y|a\in\{a_{1},a_{2},a_{3}\}\right]$. For a given question, this combines the accuracy of the final response $\hat{Y}$ with the accuracy of the current response $\hat{O}$ (if there have been re-queries), with a scaling factor $\lambda$ between the two terms. We introduced the second term because without it, we observed that if TREACLE repeatedly chose action $a_{2}$ (re-querying), this would result in multiple state transitions with 0 reward, until the final response was returned. In other words, including the second term avoids the issue of sparse rewards that resulted from the first term alone. Note that the correct response $Y$ is known only when training TREACLE; during test, the policy executes using the expected reward calculated by the trained policy.

### v2

@@ 1 Introduction
Figure 1: ${\sf\small TREACLE}$ chooses LLMs to achieve high accuracy and $\sim$85cost reduction, compared to individual LLMs.
The success of large language models (LLMs) in recent years has led to a explosion of heterogeneous models and providers, including as Meta’s Llama, OpenAI’s ChatGPT, and Google’s Gemini. As LLMs continue to proliferate in the near future, we envisage a generative AI marketplace with a large variety of providers, LLMs, and deployments. Notably, LLMs have widely varying capabilities and costs: capabilities in terms of accuracy in responding to different types of queries, and cost in terms of monetary price and query latency. As an illustration, the accuracy versus cost tradeoffs of various Llama and GPT LLMs are shown in Figure 1 on grade school math word problems [4]. As can be seen, GPT-3.5 tends to have lower accuracy than GPT-4 (79% vs 92% respectively), but costs about 20 times less. This heterogeneous array of LLMs can bewilder users who must choose between them.
Another challenge is that the specific prompt included in the question plays a critical role in eliciting accurate responses. This is especially true for reasoning problems where prompting a model to explain its reasoning can produce more accurate, but often more costly, answers. Chain-of-thought (CoT) [17] is an example of such a prompting scheme, in which the question includes a few examples of worked out problems, which cost more (due to the additional words included in the question) but also produce more accurate responses. For example, in Figure 1, GPT-4 with CoT (pink triangle) achieves a 92% accuracy, compared to GPT-4 with a domain expert prompt (brown dot, reminding the LLM that it is a “math solver”) that achieves 83%. However, using the CoT prompt costs 3.9$\times$ more due to the extra words included in the query. A final challenge is that the optimal choice of LLM and prompt depends on the specific question being asked; the accuracy of a particular LLM and prompt combination for a particular question is unknown in advance, requiring learning or prediction.
@@
Several recent works utilize multiple LLMs during inference with a cascade design, where queries propagate through a cascade of LLMs, considering the LLMs’ accuracy-cost tradeoffs. Most aim to maximize accuracy and lack an explicit way to control long-term costs, as TREACLE has. By posing the problem of LLM and prompt selection as a budget-constrained policy optimization, TREACLE provides a unified approach to efficient LLM cascades (see Table 1). TREACLE makes informed decisions based on the full context of the LLM cascade, including the query embedding, answer statistics, and remaining budget. Overall, this paper makes the following contributions:
  • $\bullet$
@@
  • $\bullet$
    An adaptive LLM and prompt selection policy based on reinforcement learning. ${\sf\small TREACLE}$ dynamically chooses the right LLM and prompt for each question. It does this by leveraging context about the current question, re-querying the models if needed to verify the consistency of the responses, and thinking ahead about the remaining budget. We also provide some theoretical justification for ${\sf\small TREACLE}$’s key design choices.
  • $\bullet$
    Extensive evaluations. We show that TREACLE substantially saves on cost while maintaining high accuracy on mathematical and commonsense reasoning tasks. We demonstrate its robustness to different budgets, question difficulty, price changes, new LLMs, and new unseen task types.
@@ 2 Related Work
FrugalGPT [3] is perhaps the closest to this work, as they considered a similar cost-constrained LLM selection problem with a threshold-based policy to select from a sorted list of LLMs. Our approach differs in several key aspects: we utilize a reinforcement learning policy that chooses both LLMs and prompts, rather than a threshold-based scheme; we utilize the full context of the current question to make decisions, including the text embedding of the current question and the history of past responses; and our method can re-query the same LLM and aggregate previous responses to estimate the correctness of the current response.
Table 1: Comparison to related works.
Query embedding Response consistency Prompt and LLM selection Long-term budget Robust to new models
FrugalGPT [3]
AutoMix [7]
MoT [20]
TREACLE
Mixture of Thought [20] explored the idea of response consistency in order to choose the right LLMs. The intuition is that higher consistency in the re-queries implies higher confidence in the correctness of the response. TREACLE employs response consistency as an input feature, along with other features, for LLM selection. AutoMix [7] introduces a “meta-verifier” to estimate whether a response is correct or a more powerful LLM is needed. Both works measure cost as a by-product of combining multiple LLMs rather than long-term constraint across all questions, as we do. Other lines of work include uncertainty estimation or prompt engineering to improve accuracy [6, 18, 20, 13, 1, 9], which is complementary to our work. The related work is summarized in Table 1.
@@ 3 Problem Statement
We study the natural language query problem of providing correct responses to a series of questions. We focus on reasoning problems (e.g., grade school math problems) because they are challenging with multiple logical steps required to reach a final correct response. The problem involves answering a sequence of $n$ questions $\mathcal{Q}$ with correct responses $\mathcal{Y}$; in other words, we have a set of questions and responses $\{(Q_{1},Y_{1}),(Q_{2},Y_{2}),\ldots,(Q_{n},Y_{n})\}$. We have a set $\mathcal{M}$ of language models (LLMs) at our disposal, which can be accessed either locally or remotely through APIs: =$\mathcal{M}=\{M_{1},M_{2},\ldots,M_{m}\}$. Also, we have a choice between $p$ prompt types, $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{p}\}$. These models and prompts have different costs (in terms of latency and monetary price) and accuracy. Each “question” can be asked multiple times to the same or different LLMs, which we call a “re-query”, in order to possibly obtain a more final accurate response.
The goal is to ensure that as many responses as possible are correct, while simultaneously minimizing the associated costs. This problem can be formulated as a Constrained Markov Decision Process (CMDP), which is represented by a tuple $(\mathcal{Q},\mathcal{S},\mathcal{A},T,r,c,\gamma,B)$, where $\mathcal{Q}$ is the ordered question set; $\mathcal{S}$ is the state space; $T:(\mathcal{Q},\mathcal{S})\times\mathcal{A}\times(\mathcal{Q},\mathcal{S})\rightarrow[0,1]$ is the transition probability function, i.e., $(Q,s)_{t+1}\sim T(\cdot|(Q,s)_{t},a_{t})$; $r:(\mathcal{Q},\mathcal{S})\times\mathcal{A}\rightarrow[R_{\text{min}},R_{\text{max}}]$ and $c:(\mathcal{Q},\mathcal{S})\rightarrow\mathbb{R}^{+}$ denote the reward and cost function; $\gamma\in[0,1]$ is the discount factor for future reward and cost; and $B$ is the total budget. A policy $\pi:(\mathcal{Q},\mathcal{S})\rightarrow P(\mathcal{A})$ maps the question-state pairs to a probability distribution over actions. A trajectory $\tau$ is composed of a sequence of (question, state)-action pairs: $\tau=\left\{\tau_{(Q,s)_{0}},\tau_{a_{0}},\tau_{(Q,s)_{1}},\tau_{a_{1}},...,\tau_{(Q,s)_{L}},\tau_{a_{L}}\right\}$, where $L$ is the total number of times the LLMs are queried. Note that $L\geq n$, due to the possible re-queries. The cumulative reward and cumulative cost of trajectory $\tau$ are denoted as $R(\tau)=\sum_{t=0}^{L}\gamma^{t}r(\tau_{(Q,s)_{t}},\tau_{a_{t}})$ and $C(\tau)=\sum_{t=0}^{L}\gamma^{t}c(\tau_{(Q,s)_{t}},\tau_{a_{t}})$, respectively. The goal of our problem is to learn a policy $\pi$ from $\mathcal{D}$ that maximizes the expected cumulative reward, while satisfying the cumulative cost at the trajectory level:
$\max_{\pi}\mathbb{E}_{\tau\sim\pi,T}[R(\tau)],\;\text{s.t.}\;\forall\tau\sim\pi,T\quad C(\tau)\leq B.$ (1)
where $\tau\sim\pi,T$ denotes that $\tau$ is generated by executing $\pi$ in $T$. By grouping the rewards and costs by question instead of enumerating all re-queries, the cumulative reward and cost can be re-written as $R(\tau)=\sum_{i=1}^{n}\texttt{reward}(Y_{i},\hat{Y}_{i})$ and $C(\tau)=\sum_{i=1}^{n}\texttt{cost}(Q_{i})$, where $\hat{Y}_{i}$ is the final response for question $Q_{i}$, $\texttt{reward}(\cdot)$ is the function that measures the correctness of the final response, $\texttt{cost}(\cdot)$ is the cost function of giving a final response $\hat{Y}_{i}$ for question $Q_{i}$.
Figure 2: Overview of TREACLE framework. TREACLE decides on the next (LLM, prompt) to query in a context-aware fashion, summarized in the state variable. It can adapt to unseen tasks by projecting the new queries into the text embedding space.
@@ 4 Proposed Framework: TREACLE
We propose the TREACLE framework, depicted in Figure 2. Let the possible unique combinations of language models and prompts be denoted by $\{(M+P)_{1},(M+P)_{2},\ldots,(M+P)_{K}\}$, where $K\leq mp$. When a new question $Q_{i}$ arrives, TREACLE starts by selecting a model $M\in\mathcal{M}$ and choosing an associated prompt $P\in\mathcal{P}$ to generate a response, denoted as $\hat{O}_{i}=M(P(Q_{i}))$. TREACLE returns this as the final response $\hat{Y}_{i}$ for this question if it has a high degree of confidence in its correctness, and deducts the cost of the question and its response from the total budget $B$. Otherwise, TREACLE can select another LLM $M$ and prompt $P$ (whose choice may be informed by the result of all previously chosen models, prompts, and their responses) and re-query. This iterative process continues until TREACLE returns a final response (based on its learned policy). TREACLE then proceeds to the next question with the remaining budget and repeats the process, until all questions have been answered or there is no remaining budget. We model the problem as a Markov decision process as described in Section 3.
@@
  • $\bullet$
    Response consistency: Records all previous responses and the normalized frequency of their occurrences. The intuition is that the consistency of the previous responses can be used as a measure of confidence in the response correctness [16, 7].
  • $\bullet$
@@
  • $\bullet$
    Current question’s text embedding: Intuitively, we want to capture the question type or difficulty, which can impact the model and prompt selection decision. TREACLE does this using a text embedding of the query [5].
  • $\bullet$
@@
  • $\bullet$
    Normalized remaining budget: Based on the remaining budget, we compute the estimated number of queries for each model prompt pair as follows: $\mathcal{B}_{k}=\frac{\text{total remaining budget}}{(\text{\# questions remaining})(\text{avg cost per query of $(M+P)_{k}$})}$. The average cost per query is estimated based on the questions seen so far. If there is a large remaining budget, TREACLE may consider re-querying with large models.
@@
By allowing re-querying (action $a_{2}$), the current action influences the next state, by impacting the question under consideration and thus the relevant state features, making this a non-trivial MDP. For $a_{3}$, we constrained the set of possible model-prompt pairs to a sorted list. In other words, instead of allowing TREACLE to select any possible model and prompting scheme, we sort the $(M+P)_{k}$ in ascending order of accuracy to cost ratio and only allow TREACLE to select the next element in this list $(M+P)_{k+1}$. The ordering is based on Proposition 1 (discussed below),
@@
Rewards. The reward function assigns a positive reward to correct responses. Specifically, $r_{\tau_{a}}\left(\tau_{(Q,s)},\tau_{(Q^{\prime},s^{\prime})}\right)=\mathbb{P}\left[\hat{Y}=Y|\tau_{a}=a_{1}\right]+\lambda\mathbb{P}\left[\hat{O}=Y|\tau_{a}\in\{a_{1},a_{2},a_{3}\}\right]$. For a given question, this combines the accuracy of the final response $\hat{Y}$ with the accuracy of the current response $\hat{O}$ (if there have been re-queries), with a scaling factor $\lambda$ between the two terms. We introduced the second term because without it, we observed that if TREACLE repeatedly chose action $a_{2}$ (re-querying), this would result in multiple state transitions with 0 reward, until the final response was returned. In other words, including the second term avoids the issue of sparse rewards that resulted from the first term alone. Note that the correct response $Y$ is known only when training TREACLE; during test, the policy executes using the expected reward calculated by the trained policy.
Design choices and justifications. We next discuss two key design choices of TREACLE and their theoretical motivation. Proofs are in Appendix B.
(1) How should the LLMs and prompts be ordered in the cascade? Recall that action $a_{3}$ moves to the next $(M+P)$ in the cascade. What is the best ordering of $(M+P)$? Consider the following simplified setting. Suppose each of the $(M+P)$ have a probability of correct response $p_{k}$ and cost $c_{k}$. If we had access to an oracle that could tell us when the response of a particular $(M+P)_{k}$ is incorrect, we could then move on and try the same question with the next option $(M+P)_{k+1}$ in the cascade. We could achieve the highest accuracy using the oracle, and would only have to worry about minimizing the cost to avoid exceeding the budget. In this setting, Proposition 1 below states that the best ordering of the $(M+P)$ options is according to the ratio $\frac{p_{k}}{c_{k}}$.
Proposition 1.
With $K$ (LLM, prompt) options, each with probability of correct answer $p_{k}$ and cost $c_{k}$, ordering the options according to their cost-normalized accuracies $\frac{p_{k}}{c_{k}}$ minimizes the total cost.
This Proposition motivates TREACLE’s ordering of the $(M+P)$ options in the cascade according to their accuracy and cost ratio. This is intuitive: instead of placing the most accurate (LLM, prompt) option early in the cascade, which might incur large cost, we first query LLMs that have high accuracy per unit cost. Note that although the setup of Proposition 1 differs from Equation 1, as the cost is the objective rather than a constraint, the trajectory resulting from the ordering in Proposition 1 is also a solution to Equation 1.
(2) Do policies that consider response consistency perform well? Recall that “response consistency” is one of the features in the state vector. We seek to understand the performance of policies that consider this feature; a simple such policy is described in Definition 1 below. It returns a final response to a question if the same response value is repeated $w$ times.
Definition 1.
For each question $Q_{i}$, an $w$-consistent policy ($w\geq 2$) sets the final response $\hat{Y}_{i}=\hat{O}_{i}$ as soon as $\exists\;\hat{O}_{i}:\mathtt{count}(\hat{O}_{i})=w$. If no such $\hat{O}_{i}$ exists, fall back to $w-1,w-2$, etc.
Definition 2 below characterizes how likely the $(M+P)$ are to return an incorrect response. A question can be asked $\Omega$ times to the $(M+P)$ options in the cascade, which may not be unique due to the re-queries.
Definition 2.
Denote the $\Omega$ LLM-prompt options by $(M+P)_{j=1}^{\Omega}$. Let $\mathbb{P}(M_{j}(P_{j}(Q_{i})))$ be the output distribution of $(M+P)_{j}$ on problem $Q_{i}$. Let $\epsilon:=\sum_{i=1}^{n}\sup_{1\leq j\leq\Omega}\sum_{\hat{O}\neq Y_{i}}\mathbb{P}(M_{j}(P_{j}(Q_{i}))=\hat{O})^{2}.$
With the definitions in hand, we can now lower bound the performance of a 2-consistent policy compared to the optimal learned algorithm in Proposition 2 below. Without loss of generality, we study the case when the reward function is the accuracy.
Proposition 2.
For the problem stated in Equation 1 that achieves $C_{*}$, the optimal expected accuracy subject to budget constraints, there exists a 2-consistent policy that achieves an accuracy of at least $C_{*}-\frac{1}{2}\Omega^{2}\epsilon$.
In other words, even a simple policy that allows for re-querying and considers response consistency can achieve close to the optimal reward. This motivates TREACLE inclusion of “response consistency” as a state feature. The proposition applies generally and allows for budgets or text embeddings, and does not require the $(M+P)$ to return accurate responses. Experimentally, we find that our learned RL policy is similar to a 2-consistent policy, as 93.02% of the responses are 2-consistent (GSM8K dataset, \$0.30 budget, $\alpha=\frac{1}{20}$). This suggests that our learned policy may not be far from optimal.